OpenStudy (anonymous):

Are 2x+4/3x+6 and 2/3 equivalent expressions? Using complete sentences explain why or why not. Please help this is my last question and I'm stuck. Thanks!

OpenStudy (anonymous):

\[\frac{ 2x + 4 }{ 3x + 6 } = \frac{ 2(x+2) }{ 3(x+2) } = \frac{ 2 }{ 3 }\] not sure how to explain it lol. been a while :P

OpenStudy (anonymous):

That right there is perfectly fine. Thank you! <3

OpenStudy (anonymous):

glad i could help :)